### v3

@@
In an important contribution, Athey, 2002 establishes conditions under which the comparative statics results in this setting exhibit monotonicity. In particular, if the DM’s utility function satisfies the single crossing property (SCP) and the information structure is ordered according to the maximum likelihood ratio (MLR), then the DM’s choice exhibits monotone comparative statics [^3] Essentially, single crossing property states that if higher action is preferred in a given state, then it must also be preferred at a higher state. If an information structure is MLR-ordered, then under higher signal realizations higher states are more likely. (MCS) - that is, higher signal realizations induce higher optimal actions.
@@
The single crossing property is an ordinal generalization of the increasing differences property [^4] A function $u$ exhibits increasing differences in $(x,\theta)$ if, $x^{\prime}>x$ and $\theta^{\prime}>\theta$ imply $u(x^{\prime},\theta^{\prime})-u(x,\theta^{\prime})>u(x^{\prime},\theta)-u(x,\theta)$. Clearly, this is a cardinal property. See Topkis, 1998.. Since it implies monotone comparative statics, and due to its straightforward economic interpretation resting on strategic complementarity, the SCP has turned out to be an extremely important tool for applications - both in applied theory as well as in the empirical literature. For instance, in mechanism design, single crossing property allows us to pin down optimal transfers. In applied work, an intimately related property has been used to solve certain optimal taxation problems (see Mirrlees, 1971). In IO entry models, single crossing property is often invoked and facilitates equilibrium analysis. In labor economics, the single crossing property is closely related to the so called Le Chatelier principle which helps characterize comparative static results (see Dekel, Quah and Sinander, 2022 for a recent work). The single crossing property also has some nice implications for preference aggregation, which proves to be quite useful in applied work (see, for instance Gans and Smart, 1996).
@@
In this paper we ask, what restrictions does the single crossing property impose on certain kinds of observable data? Often, available data - especially experimental data - may take a stochastic form. The analyst may observe only some distribution(s) over available actions and may not possess any knowledge of either the DM’s utility function or her private information. It is therefore important to understand whether the behavior of a Bayes rational agent imposes any restrictions on the stochastic data observed by the analyst. Caplin and Martin, 2015 show that rationalizability by a Bayes rational agent with access to some private information is equivalent to a single condition. Here, we ask what if the single crossing and MLR-ordered information structure impose any additional restrictions on top of those identified by Caplin and Martin, 2015. The importance of this question is illustrated by the aforementioned examples. Since the SCP has proven to be such an important property, it seems worthwhile to see if it imposes meaningful testable restrictions on observable data.
@@
Athey’s results suggest that the stochastic data should exhibit certain kind of monotonicity. If $(i)$ the DM’s utility function is single crossing; and $(ii)$ private information structure is MLR-ordered, then we know from Athey, 2002 that $(iii)$ the comparative statics are monotone. How does (i)(iii)(i)-(iii) affect the data? In our setup, neither the utility function nor the private information structure is observed. Only a set of conditional distributions over actions are observed. If $(ii)$ holds, then higher states lead to higher signal realizations more often - or with higher likelihood. If $(i)$ holds as well, then $(iii)$ is true, and so higher signals would lead directly to higher actions. One can expect this to be exhibited in data: since higher signals are more likely, and they lead to higher actions, we would expect higher actions to be more likely at higher states. This monotonicity translates into the data itself being MLR-ordered - that is, higher actions have higher relative probabilities at higher states. Proposition 13 shows that this intuition holds for the binary case. In fact, we establish something much stronger for the binary case. We show that this condition is not only both necessary and sufficient for rationalization by an SCP utility function, but such data can only be rationalized by a utility function satisfying the single crossing property. Furthermore, the main result of the paper shows that for any finite number of states, the MLR condition is both necessary and sufficient for rationalizability by a single crossing utility and an MLR ordered information structure.
@@
Related literature. Our paper tries to answer a question similar to ones asked within the literature on revealed preference theory, although the dataset we consider is quite different from finite data typically considered in classical revealed preference theory. Green and Osband, 1991, Green and Srivastava, 1986, Echenique and Saito, 2015, Bayer et al., 2013 are prominent examples of classical revealed preference theory for expected utility. However, this set of papers considers a much richer setting. They focus primarily on consumption or asset demand, while none explicitly feature information processing or consider the single crossing property. Caplin and Martin, 2015, Doval and Eilat, 2023, Rehbeck, 2023 consider stochastic datasets and consider rationalizability by a Bayesian DM. The latter two consider rationalizability when the analyst only observes marginal distribution(s) over actions, but knows the DM’s utility function. Our paper can thus be thought of as imposing particular shape restrictions in the setting of Caplin and Martin, 2015. Lazzati, Quah and Shirai, 2018 consider similar shape restrictions, but they prove a rationalizability theorem with finite data, where the variation in data comes from “budgets” of actions. With finite data, they show that single crossing is indistinguishable from a weaker property developed by Quah and Strulovici, 2009 - called interval dominance property. They also derive similar results for cross-sectional datasets and Bayesian games.
Apesteguia, Ballester and Lu, 2017 posits a single crossing random utility model, dubbed SCRUM. The primary difference between that paper and ours is contextual: they deal with stochastic choice - there is a set of preferences with a probability measure defined on them that needs to be recovered. Here, the preferences are not stochastic - only the data is - and one is interested in the existence of some preference and information structure consistent with the given data. Put another way, while the structure of the dataset available to the analyst is somewhat similar across these two groups of papers, the postulated data generating process is different.
@@ 1.1 Preliminaries
Each signal realization $s\in S$ induces a posterior distribution $\mu(\cdot|s)\in\Delta(\Theta)$. That is, if the signal realization observed by the DM is $s\in S$, then updating the prior belief using Bayes rule leads to the posterior distribution $\mu(\cdot|s)\in\Delta(\Theta)$. Thus, ex-ante, the probability of distribution $\mu(\cdot|s)$ being induced is simply the probability of signal $s$ being observed, that is $\sum_{\theta\in\Theta}\mu_{0}(\theta)\mu(s|\theta)$. It is well known, at least since Blackwell, 1951, that there is a bijection between an information structure and the distribution over posteriors induced by it. Therefore, the information structure can equivalently be represented as a mapping $\pi:\Theta\rightarrow\Delta(\Delta(\Theta))$. Going forward, both these representations shall be employed interchangeably. When the latter representation is employed, $\pi(\gamma|\theta)$ will be used to denote the probability placed by the information structure on posterior belief $\gamma$ in state $\theta$.
@@
Athey, 2002 provides conditions on $u$ and $\mathcal{I}$ under which monotone comparative statics (MCS) are obtained in this model. Some additional definitions are needed before Athey’s result can be stated.
@@
We call a Bayesian DM with a single crossing utility function and an MLR-ordered information structure a Monontone Bayesian Expected Utility Decision Maker (MBEU-DM). Among other things, Athey, 2002 shows that together, the two properties of SCP utility and MLR-ordered information imply monotone comparative statics. One version of the result is stated below.
@@
(Athey, 2002, Theorem 2) Let $\Theta,A$ be partially ordered sets. Consider a DM characterized by $u$ with access to some MLR-ordered information structure $\mathcal{I}$. If $u$ satisfies the single crossing property, then the DM exhibits monotone comparative statics. [^10] In fact, Athey, 2002 shows something much stronger: $(i)$ MCS obtains for every single crossing utility if and only if the information structure is MLR-ordered and $(ii)$ MCS obtains for every MLR-ordered information structure if and only if the utility is single crossing. That is, together, the two conditions form a minimal pair of sufficient conditions for MCS - neither of the two can be weakened any further. However, it must be noted that in the class of models considered in Athey, 2002, the set of available actions is not fixed, but changes with the parameter (in accordance with the strong set order). If the constraint set is taken to be fixed, Quah and Strulovici, 2009 show that single crossing can be further weakened to a property they call interval dominance.
@@ 1.2 Data and rationalizability
In what follows, it will be assumed that the analyst knows the set of possible states $\Theta$, the set of actions, $A$, the prior $\mu_{0}$, and stochastic choice data $\bm{q}$. Thus, the tuple $\mathcal{D}\equiv\langle A,\Theta,\mu_{0},\bm{q}\rangle$ denotes the observables available to the analyst. [^12] It is also being assumed that the (exogenously given) orders over the sets $A,\Theta$, and $S$ are known to the analyst. This corresponds to the setting considered by Caplin and Martin, 2015. There it is shown that the existence of a rationalizing utility function and an information structure is equivalent to a single condition on $\mathcal{D}$. Our paper investigates whether the single crossing property imposes further conditions on $\mathcal{D}$.
@@ 5.1 Constructing a single crossing utility function
In order to prove the result, we rely on Theorem 1 of Caplin and Martin, 2015. They show that rationalizability by a Bayesian expected utility maximizer requires the existence of a solution to a certain system of inequalities. The solution to this system is in turn a rationalizing utility function. Since the objective here is to prove the existence of a rationalizing utility that is also single crossing, it is sufficient to show that, when the dataset is MLR ordered, there exists a single crossing utility that solves the aforementioned system of inequalities.
We proceed to construct a SC utility function inductively. For each consecutive pair $a+1,a$, we compute single crossing differences $\{A^{\theta}_{a+1,a}\}_{\theta\in\Theta}$ and then find a utility function with those differences. However, an additional difficulty that needs to be addressed when $N>3$ concerns aggregation. For instance, even if $\{A^{\theta}_{a+1,a}\}$ and $\{A^{\theta}_{a,a-1}\}$ are single crossing, $\{A^{\theta}_{a+1,a-1}\}=\{A^{\theta}_{a+1,a}+A^{\theta}_{a,a-1}\}$ may not be so. That is, the single crossing property is not necessarily preserved under aggregation. However, Quah and Strulovici, 2012 report the condition - called the signed ratio property - under which the single crossing property is preserved under aggregation. These can be written as linear restrictions and incorporated into the systems of equations that characterize the single crossing differences.
@@
Exactly one of the following systems has a solution [^17] There are many different versions, and proofs of the Theorem of Alternative, sometimes also known as Farkas’ Lemma. See for instance Gale, 1989.:
@@
For l=k1,k2l=k-1,k-2, SCP follows by construction. For l<k2l<k-2, and by Claim 2, we only need to take care of cases with $A^{\theta}_{k,k-1}<0,A^{\theta}_{k-1,l}>0$. These are exactly the states $\theta$ that lie in $\widetilde{\Theta}^{k,l}$ and the restrictions involving $\bm{\kappa_{\widetilde{A}_{k}}}$ ensure that the appropriate version of signed ratio property holds for all such action-state pairs. SCP then follows by Proposition 1 of Quah and Strulovici, 2012. ∎
We now show that the differences are “optimal”. The optimality conditions are precisely the NIAS inequalities from Caplin and Martin, 2015, which can be rewritten in terms of utility differences.
@@
Once again, it follows from the Theorem 1 of Caplin and Martin, 2015 that this utility function rationalizes the data. However, we still need a choice rule and an MLR ordered information structure, that together with the utility function rationalize the data.
@@ 5.2 Constructing the information structure and the choice rule
The procedure for constructing the information structure is exactly as that found in Caplin and Martin, 2015. A key observation, however, is that when the data is MLR-ordered, then so is the rationalizing information structure. That $\langle u,\mathcal{I},C\rangle$ rationalize that data - and hence satisfy conditions 2, 3, and 4 of Definition 4 - follow directly from Theorem 1 of Caplin and Martin, 2015 (see Theorem 5). Since we have shown that $u$ is single crossing and $\mathcal{I}$ is MLR ordered, it follows that the tuple $\langle u,\mathcal{I},C\rangle$ represents an MBEU-DM that rationalizes the data.
@@ 6 Revealed Informedness
For single crossing utilities, we typically consider the Lehmann order (Lehmann, 1988), which is weaker than the Blackwell order. In particular, if an information structure $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in Blackwell order, then $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in Lehmann order, but the converse may ot hold. In particular, $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in the Lehmann order if and only if for every single-crossing utility function ex-ante utility is higher under $\mathcal{I}_{1}$ than under $\mathcal{I}_{2}$ (see for example, Persico, 2000, Jewitt, 2006). Below we give a formal definition of Lehmann order.
@@
Let $\mathcal{I}_{i}\equiv\{\mu^{i}(\cdot|\theta)\}_{\theta\in\Theta}\subset\Delta(S^{i}),i=1,2$ be MLR-ordered information structures. Let $\{F^{i}_{\theta}(\cdot)\}_{\theta\in\Theta},i=1,2$ be the corresponding distribution functions. Then, $\mathcal{I}_{1}$ is said to be Lehmann higher, or more accurate, than $\mathcal{I}_{2}$ if there exists an function $h:S^{2}\times\Theta\rightarrow S^{1}$ which is increasing in $\theta$ and, such that, for every $\theta,s^{1}\in S^{1}$, [^21] While this definition is appropriate only for continuous distributions $F^{i}$, Lehmann, 1988 shows that it is without loss of generality to assume that $F^{i}$ are continuous. That is, if $F^{1},F^{2}$ are discontinuous, there exist continuous distributions $F^{i^{*}}$ which are informationally equivalent to $F^{i}$.
@@
Suppose $\mathcal{I}_{1}=\{\mu^{1}(\cdot|s)\}_{s\in S^{1}},\mathcal{I}_{2}\equiv\{\mu^{2}(\cdot|s)\}_{s\in S^{2}}$ are MLR ordered, and let $\nu^{1},\nu^{2}$ be the corresponding marginal distributions over signals. [^22] So, $\nu^{i}(s)=\int_{\theta\in\Theta}\mu^{i}(s|\theta)\mu_{0}(\theta)$ for each $s\in S^{i}$. $\mathcal{I}_{1}$ is Lehmann higher than $\mathcal{I}_{2}$ if and only if, for every single crossing utility function $u(a,\theta)$, [^23] See Lehmann, 1988, theorem 5.1.
@@
While the proof is entirely trivial, we contend that this result is nonetheless interesting. To underscore the argument, it is instructive to compare our rationalizability result with that of Caplin and Martin (2015). Indeed, if one is interested in BEU-rationalizability alone, and the analyst has datasets $\mathcal{D}^{1},\mathcal{D}^{2}$ such that $(i)$ both $\mathcal{D}^{1},\mathcal{D}^{2}$ are BEU-rationalizable and $(ii)$ $\mathcal{D}^{1}$ Blackwell dominates $\mathcal{D}^{2}$, then the analyst can conclude that the BEU-DM corresponding to $\mathcal{D}^{1}$ is more informed than the BEU-DM corresponding to $\mathcal{D}^{2}$. But, as pointed out above, it is not known whether, in general, a solution to the NIAS inequalities exists. Secondly, Blackwell order is quite incomplete. On the other hand, the definition of Lehmann order requires the datasets to be MLR ordered - which guarantees the existence of an MBEU-DM.

### v4

@@
In an important contribution, athey2002monotone establishes conditions under which the comparative statics results in this setting exhibit monotonicity. In particular, if the DM’s utility function satisfies the single crossing property (SCP) and the information structure is ordered according to the maximum likelihood ratio (MLR), then the DM’s choice exhibits monotone comparative statics [^3] Essentially, single crossing property states that if higher action is preferred in a given state, then it must also be preferred at a higher state. If an information structure is MLR-ordered, then under higher signal realizations higher states are more likely. (MCS) - that is, higher signal realizations induce higher optimal actions.
@@
The single crossing property is an ordinal generalization of the increasing differences property [^4] A function $u$ exhibits increasing differences in $(x,\theta)$ if, $x^{\prime}>x$ and $\theta^{\prime}>\theta$ imply $u(x^{\prime},\theta^{\prime})-u(x,\theta^{\prime})>u(x^{\prime},\theta)-u(x,\theta)$. Clearly, this is a cardinal property. See topkis1998supermodularity.. Since it implies monotone comparative statics, and due to its straightforward economic interpretation resting on strategic complementarity, the SCP has turned out to be an extremely important tool for applications - both in applied theory as well as in the empirical literature. For instance, in mechanism design, single crossing property allows us to pin down optimal transfers. In applied work, an intimately related property has been used to solve certain optimal taxation problems (see mirrlees1971exploration). In IO entry models, single crossing property is often invoked and facilitates equilibrium analysis. In labor economics, the single crossing property is closely related to the so called Le Chatelier principle which helps characterize comparative static results (see dekel2022comparative for a recent work). The single crossing property also has some nice implications for preference aggregation, which proves to be quite useful in applied work (see, for instance gans1996majority).
@@
In this paper we ask, what restrictions does the single crossing property impose on certain kinds of observable data? Often, available data - especially experimental data - may take a stochastic form. The analyst may observe only some distribution(s) over available actions and may not possess any knowledge of either the DM’s utility function or her private information. It is therefore important to understand whether the behavior of a Bayes rational agent imposes any restrictions on the stochastic data observed by the analyst. caplin2015testable show that rationalizability by a Bayes rational agent with access to some private information is equivalent to a single condition. Here, we ask what if the single crossing and MLR-ordered information structure impose any additional restrictions on top of those identified by caplin2015testable. The importance of this question is illustrated by the aforementioned examples. Since the SCP has proven to be such an important property, it seems worthwhile to see if it imposes meaningful testable restrictions on observable data.
@@
Athey’s results suggest that the stochastic data should exhibit certain kind of monotonicity. If $(i)$ the DM’s utility function is single crossing; and $(ii)$ private information structure is MLR-ordered, then we know from athey2002monotone that $(iii)$ the comparative statics are monotone. How does (i)(iii)(i)-(iii) affect the data? In our setup, neither the utility function nor the private information structure is observed. Only a set of conditional distributions over actions are observed. If $(ii)$ holds, then higher states lead to higher signal realizations more often - or with higher likelihood. If $(i)$ holds as well, then $(iii)$ is true, and so higher signals would lead directly to higher actions. One can expect this to be exhibited in data: since higher signals are more likely, and they lead to higher actions, we would expect higher actions to be more likely at higher states. This monotonicity translates into the data itself being MLR-ordered - that is, higher actions have higher relative probabilities at higher states. Proposition 13 shows that this intuition holds for the binary case. In fact, we establish something much stronger for the binary case. We show that this condition is not only both necessary and sufficient for rationalization by an SCP utility function, but such data can only be rationalized by a utility function satisfying the single crossing property. Furthermore, the main result of the paper shows that for any finite number of states, the MLR condition is both necessary and sufficient for rationalizability by a single crossing utility and an MLR ordered information structure.
@@
Related literature. Our paper tries to answer a question similar to ones asked within the literature on revealed preference theory, although the dataset we consider is quite different from finite data typically considered in classical revealed preference theory. green1991revealed, green1986expected, echenique2015savage, bayer2013ambiguity are prominent examples of classical revealed preference theory for expected utility. However, this set of papers considers a much richer setting. They focus primarily on consumption or asset demand, while none explicitly feature information processing or consider the single crossing property. caplin2015testable, doval2023core, rehbeck2023revealed consider stochastic datasets and consider rationalizability by a Bayesian DM. The latter two consider rationalizability when the analyst only observes marginal distribution(s) over actions, but knows the DM’s utility function. Our paper can thus be thought of as imposing particular shape restrictions in the setting of caplin2015testable. lazzati2018nonparametric consider similar shape restrictions, but they prove a rationalizability theorem with finite data, where the variation in data comes from “budgets” of actions. With finite data, they show that single crossing is indistinguishable from a weaker property developed by quah2009comparative - called interval dominance property. They also derive similar results for cross-sectional datasets and Bayesian games.
apesteguia2017single posits a single crossing random utility model, dubbed SCRUM. The primary difference between that paper and ours is contextual: they deal with stochastic choice - there is a set of preferences with a probability measure defined on them that needs to be recovered. Here, the preferences are not stochastic - only the data is - and one is interested in the existence of some preference and information structure consistent with the given data. Put another way, while the structure of the dataset available to the analyst is somewhat similar across these two groups of papers, the postulated data generating process is different.
@@ 1.1 Preliminaries
Each signal realization $s\in S$ induces a posterior distribution $\mu(\cdot|s)\in\Delta(\Theta)$. That is, if the signal realization observed by the DM is $s\in S$, then updating the prior belief using Bayes rule leads to the posterior distribution $\mu(\cdot|s)\in\Delta(\Theta)$. Thus, ex-ante, the probability of distribution $\mu(\cdot|s)$ being induced is simply the probability of signal $s$ being observed, that is $\sum_{\theta\in\Theta}\mu_{0}(\theta)\mu(s|\theta)$. It is well known, at least since blackwell1951comparison, that there is a bijection between an information structure and the distribution over posteriors induced by it. Therefore, the information structure can equivalently be represented as a mapping $\pi:\Theta\rightarrow\Delta(\Delta(\Theta))$. Going forward, both these representations shall be employed interchangeably. When the latter representation is employed, $\pi(\gamma|\theta)$ will be used to denote the probability placed by the information structure on posterior belief $\gamma$ in state $\theta$.
@@
athey2002monotone provides conditions on $u$ and $\mathcal{I}$ under which monotone comparative statics (MCS) are obtained in this model. Some additional definitions are needed before Athey’s result can be stated.
@@
We call a Bayesian DM with a single crossing utility function and an MLR-ordered information structure a Monontone Bayesian Expected Utility Decision Maker (MBEU-DM). Among other things, athey2002monotone shows that together, the two properties of SCP utility and MLR-ordered information imply monotone comparative statics. One version of the result is stated below.
@@
(athey2002monotone, Theorem 2) Let $\Theta,A$ be partially ordered sets. Consider a DM characterized by $u$ with access to some MLR-ordered information structure $\mathcal{I}$. If $u$ satisfies the single crossing property, then the DM exhibits monotone comparative statics. [^10] In fact, athey2002monotone shows something much stronger: $(i)$ MCS obtains for every single crossing utility if and only if the information structure is MLR-ordered and $(ii)$ MCS obtains for every MLR-ordered information structure if and only if the utility is single crossing. That is, together, the two conditions form a minimal pair of sufficient conditions for MCS - neither of the two can be weakened any further. However, it must be noted that in the class of models considered in athey2002monotone, the set of available actions is not fixed, but changes with the parameter (in accordance with the strong set order). If the constraint set is taken to be fixed, quah2009comparative show that single crossing can be further weakened to a property they call interval dominance.
@@ 1.2 Data and rationalizability
In what follows, it will be assumed that the analyst knows the set of possible states $\Theta$, the set of actions, $A$, the prior $\mu_{0}$, and stochastic choice data $\bm{q}$. Thus, the tuple $\mathcal{D}\equiv\langle A,\Theta,\mu_{0},\bm{q}\rangle$ denotes the observables available to the analyst. [^12] It is also being assumed that the (exogenously given) orders over the sets $A,\Theta$, and $S$ are known to the analyst. This corresponds to the setting considered by caplin2015testable. There it is shown that the existence of a rationalizing utility function and an information structure is equivalent to a single condition on $\mathcal{D}$. Our paper investigates whether the single crossing property imposes further conditions on $\mathcal{D}$.
@@ 5.1 Constructing a single crossing utility function
In order to prove the result, we rely on Theorem 1 of caplin2015testable. They show that rationalizability by a Bayesian expected utility maximizer requires the existence of a solution to a certain system of inequalities. The solution to this system is in turn a rationalizing utility function. Since the objective here is to prove the existence of a rationalizing utility that is also single crossing, it is sufficient to show that, when the dataset is MLR ordered, there exists a single crossing utility that solves the aforementioned system of inequalities.
We proceed to construct a SC utility function inductively. For each consecutive pair $a+1,a$, we compute single crossing differences $\{A^{\theta}_{a+1,a}\}_{\theta\in\Theta}$ and then find a utility function with those differences. However, an additional difficulty that needs to be addressed when $N>3$ concerns aggregation. For instance, even if $\{A^{\theta}_{a+1,a}\}$ and $\{A^{\theta}_{a,a-1}\}$ are single crossing, $\{A^{\theta}_{a+1,a-1}\}=\{A^{\theta}_{a+1,a}+A^{\theta}_{a,a-1}\}$ may not be so. That is, the single crossing property is not necessarily preserved under aggregation. However, quah2012aggregating report the condition - called the signed ratio property - under which the single crossing property is preserved under aggregation. These can be written as linear restrictions and incorporated into the systems of equations that characterize the single crossing differences.
@@
Exactly one of the following systems has a solution [^17] There are many different versions, and proofs of the Theorem of Alternative, sometimes also known as Farkas’ Lemma. See for instance gale1989theory.:
@@
For l=k1,k2l=k-1,k-2, SCP follows by construction. For l<k2l<k-2, and by Claim 2, we only need to take care of cases with $A^{\theta}_{k,k-1}<0,A^{\theta}_{k-1,l}>0$. These are exactly the states $\theta$ that lie in $\widetilde{\Theta}^{k,l}$ and the restrictions involving $\bm{\kappa_{\widetilde{A}_{k}}}$ ensure that the appropriate version of signed ratio property holds for all such action-state pairs. SCP then follows by Proposition 1 of quah2012aggregating. ∎
We now show that the differences are “optimal”. The optimality conditions are precisely the NIAS inequalities from caplin2015testable, which can be rewritten in terms of utility differences.
@@
Once again, it follows from the Theorem 1 of caplin2015testable that this utility function rationalizes the data. However, we still need a choice rule and an MLR ordered information structure, that together with the utility function rationalize the data.
@@ 5.2 Constructing the information structure and the choice rule
The procedure for constructing the information structure is exactly as that found in caplin2015testable. A key observation, however, is that when the data is MLR-ordered, then so is the rationalizing information structure. That $\langle u,\mathcal{I},C\rangle$ rationalize that data - and hence satisfy conditions 2, 3, and 4 of Definition 4 - follow directly from Theorem 1 of caplin2015testable (see Theorem 5). Since we have shown that $u$ is single crossing and $\mathcal{I}$ is MLR ordered, it follows that the tuple $\langle u,\mathcal{I},C\rangle$ represents an MBEU-DM that rationalizes the data.
@@ 6 Revealed Informedness
For single crossing utilities, we typically consider the Lehmann order (Lehman), which is weaker than the Blackwell order. In particular, if an information structure $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in Blackwell order, then $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in Lehmann order, but the converse may ot hold. In particular, $\mathcal{I}_{1}$ is higher than $\mathcal{I}_{2}$ in the Lehmann order if and only if for every single-crossing utility function ex-ante utility is higher under $\mathcal{I}_{1}$ than under $\mathcal{I}_{2}$ (see for example, Persico, Jewitt). Below we give a formal definition of Lehmann order.
@@
Let $\mathcal{I}_{i}\equiv\{\mu^{i}(\cdot|\theta)\}_{\theta\in\Theta}\subset\Delta(S^{i}),i=1,2$ be MLR-ordered information structures. Let $\{F^{i}_{\theta}(\cdot)\}_{\theta\in\Theta},i=1,2$ be the corresponding distribution functions. Then, $\mathcal{I}_{1}$ is said to be Lehmann higher, or more accurate, than $\mathcal{I}_{2}$ if there exists an function $h:S^{2}\times\Theta\rightarrow S^{1}$ which is increasing in $\theta$ and, such that, for every $\theta,s^{1}\in S^{1}$, [^21] While this definition is appropriate only for continuous distributions $F^{i}$, Lehman shows that it is without loss of generality to assume that $F^{i}$ are continuous. That is, if $F^{1},F^{2}$ are discontinuous, there exist continuous distributions $F^{i^{*}}$ which are informationally equivalent to $F^{i}$.
@@
Suppose $\mathcal{I}_{1}=\{\mu^{1}(\cdot|s)\}_{s\in S^{1}},\mathcal{I}_{2}\equiv\{\mu^{2}(\cdot|s)\}_{s\in S^{2}}$ are MLR ordered, and let $\nu^{1},\nu^{2}$ be the corresponding marginal distributions over signals. [^22] So, $\nu^{i}(s)=\int_{\theta\in\Theta}\mu^{i}(s|\theta)\mu_{0}(\theta)$ for each $s\in S^{i}$. $\mathcal{I}_{1}$ is Lehmann higher than $\mathcal{I}_{2}$ if and only if, for every single crossing utility function $u(a,\theta)$, [^23] See Lehman, theorem 5.1.
@@
While the proof is entirely trivial, we contend that this result is nonetheless interesting. To underscore the argument, it is instructive to compare our rationalizability result with that of caplin2015testable. Indeed, if one is interested in BEU-rationalizability alone, and the analyst has datasets $\mathcal{D}^{1},\mathcal{D}^{2}$ such that $(i)$ both $\mathcal{D}^{1},\mathcal{D}^{2}$ are BEU-rationalizable and $(ii)$ $\mathcal{D}^{1}$ Blackwell dominates $\mathcal{D}^{2}$, then the analyst can conclude that the BEU-DM corresponding to $\mathcal{D}^{1}$ is more informed than the BEU-DM corresponding to $\mathcal{D}^{2}$. But, as pointed out above, it is not known whether, in general, a solution to the NIAS inequalities exists. Secondly, Blackwell order is quite incomplete. On the other hand, the definition of Lehmann order requires the datasets to be MLR ordered - which guarantees the existence of an MBEU-DM.